FIND EQUATION OF THE LINE PASSING THROUGH THE POINTS (-2,2) AND (8,-3)
Answer by fractalier(6550)

No need to YELL.
First find the slope, m.
m = (-3-2)/(8-(-2)) = -5/10 = -1/2
Now plug into y=mx+b and find b...use the first point's coordinates...
2 = (-1/2)(-2) + b
2 = 1 + b
b = 1 so that
y = (-1/2)x + 1 is your line.